OpenStudy (anonymous):

Find mass of excess reactant. 3Cu+8HNO3-->3CU(NO3)2+2NO+4H2O. I found the mass of H2O which would be 1.5g. But I am not getting the right mass of the excess

OpenStudy (anonymous):

PLEASE SOMEONE HELP GOD WILL BLESS YOU

OpenStudy (zale101):

@nincompoop

OpenStudy (anonymous):

11g of Cu and HNO3 were used !

OpenStudy (anonymous):

ugh :(

OpenStudy (anonymous):

okai so you have your equation Find which is limiting and which is excess

OpenStudy (anonymous):

like in mol

OpenStudy (anonymous):

Limiting is Cu with 0.173 moles. and Excess is HNO3 which is 0.174 moles

OpenStudy (anonymous):

(limiting would be the smaller number right?)

OpenStudy (anonymous):

yes it's the smaller number lemme check that

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

they don't react in 1:1 ratio--it's 3 Cu + 8 HNO3, so 3:8 ratio.

OpenStudy (anonymous):

so what would I do ?

OpenStudy (anonymous):

for HNO3, use 0,175mol instead

OpenStudy (anonymous):

it was 0,1745 so yeah anyways now that you found your limiting and excess....

OpenStudy (anonymous):

calculate how much excess is consumed in the reaction Excess is HNO3 Use mole ratio to determine it The equation follows: mol of limiting*(mole ratio) = mol of excess used up

OpenStudy (nincompoop):

so what is the problem?

OpenStudy (anonymous):

so 0.173moles of limiting *(3/8) which gave me 0.064 :S

OpenStudy (nincompoop):

could you show me how you are solving this?

OpenStudy (anonymous):

incorrect it's 8/3 and ill show you a sample nincompoop

OpenStudy (anonymous):

...after

OpenStudy (anonymous):

so I got 0.461

OpenStudy (nincompoop):

what sample? just show me the dimensional analysis

OpenStudy (anonymous):

alright lemme draw it out in the meantime use that mole value you got and subtract it from your original excess

OpenStudy (anonymous):

wait, nope, not gonna work...okaay so ill just draw it out

OpenStudy (anonymous):

okay so 0.4613-0.175 = 0.2863

OpenStudy (anonymous):

um okayy

OpenStudy (nincompoop):

what difficulty are you having converting moles to grams?

OpenStudy (anonymous):

well I'm not getting the right answer soo....

OpenStudy (anonymous):

I dont know what im doing wrong

OpenStudy (nincompoop):

make your life easier, write down the masses of each reactants how many grams are there in one mole of Cu?

OpenStudy (anonymous):

1. List Ratio 2. List Molar Masses 3. Find Limiting and Excess 4.Find how much mole of excess is consumed 5. Subtract from original moles 6. Convert value to desired unit

OpenStudy (nincompoop):

start using dimensional analysis

OpenStudy (anonymous):

i never really learned it.. so i made a list

OpenStudy (anonymous):

ohkaaaaaaay nin

OpenStudy (anonymous):

Okay so this is what I have so far guys. 11g of Cu so 0.173 moles. (limiting) and 11g of HNO3 so 0.174 moles. 0.174*(4H2O/8HNO3) = 0.087 x 18.0.148 which gave me 1.5g mass for H2O

OpenStudy (anonymous):

but you want mass of excess reactant

OpenStudy (anonymous):

Then I said I have 0.174moles - 0.087moles needed = 0.087 moles x 63.0119 (HNO3) which gave me 5.5g but thats wrong.

OpenStudy (anonymous):

what am i doing wrong

OpenStudy (nincompoop):

hey let me show you something that helps how do you turn a rabbit into frog? \[Rabbit \times \left( \frac{ Frog }{ Rabbit } \right)= Frog\]

OpenStudy (anonymous):

uh cancel out the rabbits o.o

OpenStudy (anonymous):

Wait can you do the question like type it out and show me step by step what you did?

OpenStudy (nincompoop):

that's what dimensional analysis is. can you show me your solution using dimensional analysis so we can pick up where you are having a problem.

OpenStudy (anonymous):

I will attach a picture and show you what I did because I didn't really learn that way

OpenStudy (nincompoop):

you have to learn that way so when you go back you can always trace where the error is.

OpenStudy (anonymous):

can you see?

OpenStudy (nincompoop):

it should look like this \[3molCu \times \left( \frac{ 63.546 gCu }{ 1molCu } \right) = \]

OpenStudy (nincompoop):

you cancel out the mol and you multiply 3 by the gram value of one mole of Cu.

OpenStudy (anonymous):

so 190.638?

OpenStudy (nincompoop):

yes, 3 moles of Cu = 190.638 you're just converting values from moles to grams, don't get excited that's not the answer yet

OpenStudy (anonymous):

lool okay so whats the next step?

OpenStudy (nincompoop):

what do you think is the next step?

OpenStudy (nincompoop):

you're identifying the excess right?

OpenStudy (anonymous):

yeahh

OpenStudy (nincompoop):

okay how would you achieve that? define excess

OpenStudy (anonymous):

excess, meaning the leftover? the extra? Soo would I subtract from something?

OpenStudy (anonymous):

the limiting or mass of h2o ?

OpenStudy (nincompoop):

\[8molHNO_3\times \left( \frac{ ?gHNO_3 }{ 1molHNO_3 } \right) = \]

OpenStudy (anonymous):

I got 504.0952

OpenStudy (nincompoop):

if I were to make a bicycle. how many wheels do I need for every frame?

OpenStudy (anonymous):

2?

OpenStudy (nincompoop):

if I were to make bicycles and I have 30 wheels and have 10 frames. how many bicycles I can make?

OpenStudy (anonymous):

I'm guessing 15? Or wait how many frames does one bike have? 2? is 2 then 20 bikes..

OpenStudy (anonymous):

NO omg I feel so stupid now

OpenStudy (nincompoop):

no… for every frame you need 2 wheels to make 1 bike 10 frames = 20 wheels this gives you 10 wheels extra

OpenStudy (anonymous):

oh..okay

OpenStudy (nincompoop):

you get this part right? how about if I had 10 wheels and 10 frames? how many bikes I can make?

OpenStudy (anonymous):

5? because you wouldn't have enough wheels?

OpenStudy (nincompoop):

awesome! in the first bicycle question, your limiting would be frames; in the second bicycle it would be the wheels. so basically your limiting is always the thing that runs out first and of course the other would be your excess.

OpenStudy (nincompoop):

is this making any sense and are you able to draw connection from bikes to chemistry?

OpenStudy (anonymous):

yess

OpenStudy (nincompoop):

do you think you can take a stab at the question once more, but this time using the dimensional analysis?

OpenStudy (anonymous):

Um. Well you said frames = 2 wheels. and frames is my limiting. So my limiting is 190.638 and my excess is 504.095. Would I divide?

OpenStudy (nincompoop):

if you're looking for the ratio then you divide

OpenStudy (anonymous):

Soooo 504.095/190.638 ?

OpenStudy (anonymous):

WAIT I GOT 2.64

OpenStudy (anonymous):

omfgg thats it?

OpenStudy (anonymous):

is taht it?

OpenStudy (anonymous):

it took my 2 hours to get to that? and all I had to do was divide the things ? wooooww. thanks tho !

OpenStudy (nincompoop):

you work your way back. this is a good example http://www.chem.tamu.edu/class/majors/tutorialnotefiles/limiting.htm

OpenStudy (nincompoop):

you're way ahead… lol I told you that you're just merely converting

OpenStudy (anonymous):

man but I tried this with another question and it didn't work. I'm getting so fusterated now

OpenStudy (anonymous):

what did I do wrong again? =l

OpenStudy (nincompoop):

lol you didn't even finish the first problem yet

OpenStudy (nincompoop):

or did you?

OpenStudy (anonymous):

wait I didnt? oh no :(

OpenStudy (nincompoop):

did you read the link I sent you? it's a good example, but I can show you how it is done from textbooks.

OpenStudy (anonymous):

after I divied 504.0942g/190.638g I got 2.64 g

OpenStudy (anonymous):

isn't that the answer?

OpenStudy (anonymous):

It's 2am, I just wanna know how to do it straight forward :/ I havea test tomorrow and I'm very nervous about it

OpenStudy (nincompoop):

check page 107 of that book

OpenStudy (anonymous):

alrighty thank you! :)

OpenStudy (anonymous):

wow this book is good, which country are you from? damn I wish we had these

OpenStudy (anonymous):

oh so theorictical is the limiting? and then they give you actual?

OpenStudy (nincompoop):

you're welcome. There's a reason why I made sure you understood the bike problem first, because it is the very same concept you will use for stoichiometry 4 wheels + 3 frames ----> 2 bikes + 1 piece of metal you identify how many wheels would it require you to make 2 bikes you do the same with frames you identify the limiting reagent or excess you convert their values to grams (or to moles if you were given in grams initially)

OpenStudy (nincompoop):

no. theoretical yield versus actual yield. theoretical is what would you produce assuming 100% efficiency, actual well of course if you did the actual experiment, how much would you produce while factoring the spills, mistakes etc.